# Entropic Trade-off Relations in Stochastic Thermodynamics via Replica Markov Processes

### v2

@@ IV Conclusion
Substituting Eq. (84) into Eq. (17), we obtain
@@
Here, $X^{(k)}(\tau)$ is the state of the $k$-th process at time $t=\tau$. Therefore, Eq. (86) is 0 if the final states of the $K$ processes agree. In this study, when considering a random walker on a network, we assume that the initial state of the random walk is $B_{\mu_{0}}$. This implies that, when there is no jump within the interval $0,\tau]$, the final state at time $t=\tau$ is $X^{(1)}(\tau)=X^{(2)}(\tau)=\cdots=X^{(K)}(\tau)=B_{\mu_{0}}$. Therefore, the observable $\overline{\Xi}^{(\otimes K)}$ satisfies
@@
Here, $\mathcal{B}(\tau)$ is the quantum dynamical activity, whose expression is expressed in Appendix F. We consider the $K$ replica scenario, where there are $K$ identical processes. Let $\mathfrak{p}^{(\otimes K)}_{Q}(\tau)$ denote the probability that there is no jump in the $K$ replica processes. Since each process is independent, from Eq. (90), $\mathfrak{p}^{(\otimes K)}_{Q}(\tau)$ satisfies
@@
$\overline{\Lambda}_{Q\circ}^{(\otimes K)}(\Gamma_{Q}^{\otimes K})\equiv\begin{cases}0&N_{Q\circ}^{(1)}(\Gamma_{Q}^{(1)})=N_{Q\circ}^{(2)}(\Gamma_{Q}^{(2)})=\cdots=N_{Q\circ}^{(K)}(\Gamma_{Q}^{(K)})\\ 1&\mathrm{otherwise}\end{cases}.$ (95)
Then $\overline{\Lambda}_{Q\circ}^{(\otimes K)}$ returns 0 only when the values of the observable $N_{Q\circ}$ in $K$ processes agree, $N_{Q\circ}^{(1)}(\Gamma_{Q}^{(1)})=N_{Q\mathrm{\circ}}^{(2)}(\Gamma_{Q}^{(2)})=\cdots=N_{Q\circ}^{(K)}(\Gamma_{Q}^{(K)})$. Then the observable $\overline{\Lambda}_{Q\circ}^{(\otimes K)}$ satisfies
@@
which is the condition of Eq. (47). Let $\mathcal{Z}_{Q\circ}$ be the set of all possible values that the observable $N_{Q\circ}(\Gamma_{Q})$ can take. Then, the expectation becomes the following:
@@
By considering a general $K$ case and substituting Eq. (97) into Eq. (52), we obtain
@@
Appendix F Quantum dynamical activity
@@
Next, we consider a quantum generalization of the dynamical activity. The quantum dynamical activity plays a central role in trade-off relations [17, 44, 65, 64, 66, 69]. The exact expression for $\mathcal{B}(t)$ in the quantum dynamical activity was recently obtained in Refs. [65, 64]. Specifically, Ref. [64] showed the following relation:

### v3

@@ IV Conclusion
The expectation and variance of $\overline{\Lambda}_{\circ}^{(\otimes K)}$ are
$\mathbb{E}[\overline{\Lambda}_{\circ}^{(\otimes K)}(\Gamma^{(\otimes K)})]$ $=1-\sum_{z\in\mathcal{Z}_{\circ}}P(N_{\circ}=z)^{K},$ (87)
$\mathrm{Var}[\overline{\Lambda}_{\circ}^{(\otimes K)}(\Gamma^{(\otimes K)})]$ $=\left(1-\sum_{z\in\mathcal{Z}_{\circ}}P(N_{\circ}=z)^{K}\right)$
$-\left(1-\sum_{z\in\mathcal{Z}_{\circ}}P(N_{\circ}=z)^{K}\right)^{2}.$ (88)
Substituting Eqs. (87) and (88) into Eq. (18), we obtain
@@
Here, $X^{(k)}(\tau)$ is the state of the $k$-th process at time $t=\tau$. Therefore, Eq. (90) is 0 if the final states of the $K$ processes agree. When deriving Eqs. (35) and (36), we assume that the initial state of the random walk is $B_{\mu_{0}}$. This implies that when there is no jump within the interval $0,\tau]$, the final state at time $t=\tau$ is $X^{(1)}(\tau)=X^{(2)}(\tau)=\cdots=X^{(K)}(\tau)=B_{\mu_{0}}$. Therefore, the observable $\overline{\Xi}^{(\otimes K)}$ satisfies
@@
Here, $\mathcal{B}(\tau)$ is the quantum dynamical activity, whose expression is expressed in Appendix G. We consider the $K$ replica scenario, where there are $K$ identical processes. Let $\mathfrak{p}^{(\otimes K)}_{Q}(\tau)$ denote the probability that there is no jump in the $K$ replica processes. Since each process is independent, from Eq. (94), $\mathfrak{p}^{(\otimes K)}_{Q}(\tau)$ satisfies
@@
$\overline{\Lambda}_{Q\circ}^{(\otimes K)}(\Gamma_{Q}^{\otimes K})$
$\equiv\begin{cases}0&N_{Q\circ}^{(1)}(\Gamma_{Q}^{(1)})=N_{Q\circ}^{(2)}(\Gamma_{Q}^{(2)})=\cdots=N_{Q\circ}^{(K)}(\Gamma_{Q}^{(K)})\\ 1&\mathrm{otherwise}\end{cases}.$ (99)
$\overline{\Lambda}_{Q\circ}^{(\otimes K)}$ returns 0 only when the values of the observable $N_{Q\circ}$ in the $K$ processes agree, $N_{Q\circ}^{(1)}(\Gamma_{Q}^{(1)})=N_{Q\mathrm{\circ}}^{(2)}(\Gamma_{Q}^{(2)})=\cdots=N_{Q\circ}^{(K)}(\Gamma_{Q}^{(K)})$. Then, the observable $\overline{\Lambda}_{Q\circ}^{(\otimes K)}$ satisfies
@@
which is the condition of Eq. (48). Let $\mathcal{Z}_{Q\circ}$ be the set of all possible values that the observable $N_{Q\circ}(\Gamma_{Q})$ can take. The expectation becomes the following:
@@
By considering the general case for $K$ and substituting Eq. (101) into Eq. (54), we obtain
@@
Appendix G Quantum dynamical activity
@@
Next, we consider a quantum generalization of the dynamical activity. The quantum dynamical activity plays a central role in trade-off relations [19, 46, 70, 69, 71, 78]. The exact expression for $\mathcal{B}(t)$ in the quantum dynamical activity was obtained in Refs. [70, 69]. Specifically, Ref. [69] showed the following relation: